# Científicas (CSIC) and Universitat de València (UV), 46980 Valencia, Spain

### v1

@@ 1 Introduction
Studying the Higgs boson more precisely at a future Higgs factory has been identified as one of the main priorities for particle physics CERN-ESU-015 ; osti_2368847 . A circular $e^{+}e^{-}$ collider is arguably the best option for refining Higgs measurements to the sub-percent level and probing the Standard Model (SM) as generally as possible across all its intricately interconnected sectors Blondel:2024mry . With an extensive physics programme consisting of runs at the $Z$ pole, $WW$, $ZH$, and $t\bar{t}$ thresholds, it would not only guarantee results that would improve tremendously our picture of the fundamental building blocks of the universe, but also enable a far wider general exploration of the TeV scale than naïvely expected. In particular, recent studies have highlighted how indirect quantum effects at the $Z$ pole, accessible by high precision measurements, can be a powerful and versatile probe of heavy new physics (NP) Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds ; Allwicher:2024sso ; Erdelyi:2024sls ; Gargalionis:2024jaw .
@@
In this paper we further study the complementarity between high accuracy precision measurements on the $Z$ pole and measurements at higher energy runs that we collectively denote as “off-pole”. It is typically thought that the $Z$ pole, while strongly constraining the oblique $\hat{S}$ and $\hat{T}$ parameters Peskin:1991sw , is not particularly sensitive to modifications of Higgs couplings, pure gauge operators, or four-fermion contact interactions that are best constrained at higher energies. The LHC has also demonstrated how energy helps accuracy when compared to LEP constraints on the $\hat{W}$ and $\hat{Y}$ parameters Farina:2016rws . However, recent work has shown that hundreds of fermionic operators in the SM Effective Field Theory (SMEFT) can be efficiently probed on-pole at FCC-ee at Next-to-Leading-Order (NLO) Allwicher:2023shc ; Dawson:2022bxd ; Bellafronte:2023amz . Here we extend applications of the “accuracy complements energy” principle by pointing out that the Higgs-only and gauge-only bosonic operators can also be constrained at the $Z$-pole with similar or better accuracy than off-pole measurements in higher energy runs, even for operators that are energy enhanced.
@@
where $\epsilon_{Z}$ and $\epsilon_{ZH}$ account for any reduction from the naive statistical expectation due to experimental and theoretical uncertainties Bernardi:2022hny ; Freitas:2019bre , and we defined an improvement factor $\Delta^{NLO/LO}$. Assuming conservatively that $\epsilon_{Z}\sim 10^{-1}$ and $\epsilon_{ZH}\sim 1$, we see for $N_{Z}\sim 10^{12}$ and $N_{ZH}\sim 10^{6}$ that indeed $\Delta_{Z/ZH}$ can be $O(1)$. We may therefore expect competitive sensitivity at NLO on the $Z$ pole with respect to off-pole constraints at leading order for any Wilson coefficient that both have access to.
This argument also applies to new physics entering at leading order in both $Z$ pole and off-pole observables but with effects that grow with energy, as for the oblique $\hat{W}$ and $\hat{Y}$ parameters Maksymyk:1993zm ; Barbieri:2004qk . An energy enhancement at higher energies, for example in $WW$ measurements, can be compensated for by higher statistical accuracy on-pole if
@@
where we defined the improvement factor $\Delta^{LO/LO}$ to include energy-enhanced terms. For $WW$ energies $E_{WW}\sim 200$ GeV, $\epsilon_{Z}\ \sim 10^{-1}$, $\epsilon_{WW}\sim 1$, $N_{Z}\sim 10^{12}$ and $N_{WW}\sim 10^{7}$, we obtain $\Delta^{LO/LO}_{Z/WW}=O(1)$. Thus, unlike at LEP (where $N_{Z}\sim 10^{7}$ and $N_{WW}\sim 10^{5}$), a Tera-Z factory is expected to be just as sensitive to the $\hat{W}$ and $\hat{Y}$ parameters on the $Z$ pole as in higher-energy $WW$ runs. This was recently shown explicitly by comparing $e^{+}e^{-}\rightarrow\bar{f}f$ off-pole processes with $Z$-pole data in Ref. Greljo:2024ytg .
Figure 1: Projected sensitivity at FCC-ee of measurements on the $Z$ pole and off-pole at higher energy runs for Higgs, gauge, and four-fermion quark/lepton operators. The colour gradient denotes the bound on the scale $\Lambda$ in TeV for unit Wilson coefficient.
Fig. 1 summarises the complementarity between on-pole and off-pole constraints. The effects of heavy new physics are parametrised by the Wilson coefficients $\mathcal{C}_{i}\equiv C_{i}/\Lambda^{2}$ of dimension-6 operators $Q_{i}$ in the Warsaw basis of the SMEFT Grzadkowski:2010es , while the projected sensitivity at FCC-ee to the effective scale of new physics is represented by the colour-coded heatmap gradient. The operators shown are grouped by Higgs, gauge, and various four-fermion operators, and are typically thought to best be constrained at leading order in the off-pole $WW$, $ZH$, and $\bar{t}t$ higher-energy runs. Darker shading indicates a stronger constraint, and we see in many cases the on-pole constraint dominating or giving a similar order of magnitude sensitivity to the corresponding off-pole constraint. This is indeed the case for the $\hat{W}$ and $\hat{Y}$ oblique parameters that map on to the $\mathcal{S}_{2W}$ and $\mathcal{S}_{2B}$ operator coefficients in the SILH basis Giudice:2007fh ; Contino:2013kra . For other operators the off-pole constraint can win due to energy enhancement effects, though the benefit of entering at leading order vanishes when considering four-fermion operators involving second and third generation fermions. Moreover, including on-pole data eliminates directions of limited sensitivity in a global fit which can remove degeneracies and improve constraints when fitting to several correlated operators at a time, as is necessary in most realistic ultraviolet (UV) models.
@@
We find this to be the case in several specific UV models that we study as examples. In the real scalar singlet and custodial quadruplet Durieux:2022hbu models, it is essential to take into account the correlated contributions to the Higgs self-coupling and Higgs self-energy when estimating the projected sensitivities in a global fit, even when there is a loop hierarchy between the two as for the weak quadruplet. Including these correlations weaken the bounds from off-pole measurements compared to considering only individual constraints; however, adding on-pole data then strengthens them further. For the real singlet scalar in particular, the on-pole information is crucial for potentially covering the entire region of parameter space in which the scalar can be a so-called “loryon” Banta:2021dek , thus closing the last remaining open window in being able to conclusively answer whether any other particles outside the SM get most of their mass from the Higgs Crawford:2024nun . We also consider the case of Weakly Interacting Massive Particles (WIMPs) and show that electroweak precision can significantly improve on the picture after HL-LHC for this elusive class of BSM extensions.
@@ 2 Setup and Observables
The measurements used in our analysis are classified as on-pole for $Z$- and $W$-pole observables, and as off-pole observables for everything else. For on-pole data, we use the following $Z$-pole observables,
@@
For these observables, we use the expressions calculated at NLO in the SMEFT with {$G_{F}$, $m_{Z}$, $\alpha$} as input parameters in Ref. Bellafronte:2023amz . The relative uncertainties and projected error reductions for the $Z$- and $W$-pole observables are given in Table 1. The $Z$ pole uncertainties are taken from Ref. DeBlas:2019qco , including the projected theoretical uncertainty only for $\Gamma_{Z}$ Bernardi:2022hny ; Freitas:2019bre . The $W$ pole uncertainties are from Ref. Bernardi:2022hny .
@@
Table 1: Projected FCC-ee improvement for $Z$-pole observables and $W$ pole observables from Bernardi:2022hny ; DeBlas:2019qco . The $A_{\ell}^{**}$ are from lepton polarization and LR asymmetry measurements at SLC. For $A_{e}^{**}$, we use the projection in deBlas:2022ofj .
On the other hand, the off-pole processes considered here are
$O_{\text{off-pole}}=\left\{\sigma\left(e^{+}e^{-}\to W^{+}W^{-}\right),\,\sigma\left(e^{+}e^{-}\to ZH\right),\,\sigma\left(e^{+}e^{-}\to f\bar{f}\right)\,,\mu_{b\bar{b}},\mu_{c\bar{c}},\mu_{\tau\bar{\tau}},\mu_{\mu\bar{\mu}}\,\right\}.$
We use the expressions in Allwicher:2023aql for the Higgs signal strengths $\mu_{f\bar{f}}$. For the $ZH$ production cross section, we use the results from Ref. Asteriadis:2024xts , which have been computed at NLO in the SMEFT. We use our own leading order (LO) calculation for the $WW$ process, given in Appendix A. For both calculations, the {$G_{F}$, $m_{W}$, $m_{Z}$} input scheme is adopted. [^2] While this is an arbitrary choice which should and does not have a significant effect in a global analysis Biekotter:2023xle , it could in principle be problematic. However, since neither $\alpha$ nor $m_{W}$ are measured precisely off-pole (above $160$ GeV), we expect that any input scheme dependence is negligible. The expected precision at which FCC-ee can measure $\sigma(ZH)$ and $\sigma(WW)$ is taken from Ref. DeBlas:2019qco .
Regarding the $e^{+}e^{-}\to f\bar{f}$ process, we use the results of the recent dedicated flavor-tagging analysis for FCC-ee Greljo:2024ytg . They use the cross-section ratio observables
@@
where the sum runs over all the light quarks. Our predictions are implemented following Appendix B of Greljo:2024ytg .
Finally, our likelihood is constructed by combining all these observables and predictions without making any flavour assumptions, with the fit performed via a simple $\chi^{2}$ procedure as in Ref. Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds . The LO $Z/W$-pole part of this likelihood has previously been applied and validated in Refs. Allwicher:2023aql ; Davighi:2023evx ; Allwicher:2023shc ; Stefanek:2024kds , and has also been cross-checked here with an independent implementation in fitmaker Ellis:2020unq .
3 Accuracy complements energy: an EFT analysis
@@
where $Q_{i}$ are dimension-6 operators and $C_{i}$ are the dimensionless Wilson coefficients generated by heavy new physics at a scale $\Lambda$. We will take $\Lambda=1$ TeV unless otherwise stated. It will be convenient to use the SILH basis Giudice:2007fh ; Contino:2013kra to characterise pure gauge operators, for which we adopt the following notation for the Wilson coefficients and operators,
$\mathcal{L}_{\rm SMEFT}\supset\sum_{i}\frac{S_{i}}{\Lambda^{2}}\mathcal{O}_{i}^{\rm SILH}\,.$ (4)
For more concise expressions we will sometimes employ dimensionful Wilson coefficients defined as $\mathcal{C}_{i}\equiv C_{i}/\Lambda^{2}$ and $\mathcal{S}_{i}\equiv S_{i}/\Lambda^{2}$.
The SMEFT language provides a systematic way of exploring modifications from heavy BSM physics that are correlated across different measurements at various energies (for some recent global fits, see e.g. Refs. Ellis:2020unq ; Bissmann:2020mfi ; Ethier:2021bye ; Grunwald:2023nli ; Garosi:2023yxg ; Allwicher:2023shc ; Bartocci:2023nvp ; Celada:2024mcf ; deBlas:2022ofj ; Bartocci:2024fmm ). Our approach is to focus on off-pole operators that enter the $Z$-pole at one higher loop order, as well as energy-enhanced off-pole operators that also enter at the $Z$-pole. We begin by illustrating our point with Higgs sector modifications before considering pure gauge operators and then summarising recent four-fermion results.
@@ 3.1 Higgs coupling modifications
Figure 4: New physics affecting the $hVV$ coupling in (a), NLO contribution of $hVV$ coupling modifications to $Z$-pole observables. Red squares indicate SMEFT contributions to the vertex.
A Higgs factory will be sensitive at leading order to new physics modifying the Higgs self-energy in the process $e^{+}e^{-}\to ZH$, as illustrated by the propagator correction in Fig. 2; we also see in Fig. 2 that it necessarily also enters at NLO in the oblique parameters on the $Z$ pole. Similarly, Fig. 3 shows that a vertex correction to the Higgs self-coupling will enter at NLO in $e^{+}e^{-}\to ZH$ and at NNLO on the $Z$ pole, while Fig. 4 shows a LO $hVV$ vertex correction modifying the $Z$ pole oblique parameter at NLO. As argued previously, we expect the suppression from an extra loop factor on the $Z$ pole to be compensated by increased statistics, such that the on-pole vs. off-pole improvement factor on constraints from all these Higgs coupling modifications could be $\Delta^{NLO/LO}_{Z/ZH}\sim\mathcal{O}(1)$ for both the ratios NLO/LO and NNLO/NLO.
@@
Figure 5: FCC-ee projected sensitivities at 68% CL to Higgs coupling modifications on and off the $Z/W$ pole. Left: SMEFT operators that modify the Higgs 2- and 3-point functions. The dashed region is obtained by linearizing Eq. 12. Right: SMEFT operators giving $hVV$ coupling modifications. The Wilson coefficients are renormalised at $\Lambda=1$ TeV.

### v2

@@ 1 Introduction
Studying the Higgs boson more precisely at a future Higgs factory has been identified as one of the main priorities for particle physics CERN-ESU-015 ; osti_2368847 . A circular $e^{+}e^{-}$ collider is arguably the best option for refining Higgs measurements to the sub-percent level and probing the Standard Model (SM) as generally as possible across all its intricately interconnected sectors Blondel:2024mry . With an extensive physics programme consisting of runs at the $Z$ pole, $WW$, ZHZH, and $t\bar{t}$ thresholds, it would not only guarantee results that would improve tremendously our picture of the fundamental building blocks of the universe, but also enable a far wider general exploration of the TeV scale than naïvely expected. In particular, recent studies have highlighted how indirect quantum effects at the $Z$ pole, accessible by high precision measurements, can be a powerful and versatile probe of heavy new physics (NP) Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds ; Allwicher:2024sso ; Erdelyi:2024sls ; Gargalionis:2024jaw .
@@
In this paper we further study the complementarity between high accuracy precision measurements on the $Z$ pole and measurements at higher energy runs that we collectively denote as “above-pole”. It is typically thought that the $Z$ pole, while strongly constraining the oblique $\hat{S}$ and $\hat{T}$ parameters Peskin:1991sw , is not particularly sensitive to modifications of Higgs couplings, pure gauge operators, or four-fermion contact interactions that are best constrained at higher energies. The LHC has also demonstrated how energy helps accuracy when compared to LEP constraints on the $\hat{W}$ and $\hat{Y}$ parameters Farina:2016rws . However, recent work has shown that hundreds of fermionic operators in the SM Effective Field Theory (SMEFT) can be efficiently probed on-pole at FCC-ee at Next-to-Leading-Order (NLO) Allwicher:2023shc ; Dawson:2022bxd ; Bellafronte:2023amz . Here we extend applications of the “accuracy complements energy” principle by pointing out that the Higgs-only and gauge-only bosonic operators can also be constrained at the $Z$-pole with similar or better accuracy than above-pole measurements in higher energy runs, even for operators that are energy enhanced.
@@
where $\epsilon_{Z}$ and $\epsilon_{ZH}$ account for any reduction from the naive statistical expectation due to experimental and theoretical uncertainties Bernardi:2022hny ; Freitas:2019bre , and we defined an improvement factor $\Delta^{NLO/LO}$. Assuming conservatively that $\epsilon_{Z}\sim 10^{-1}$ and $\epsilon_{ZH}\sim 1$, we see for $N_{Z}\sim 10^{12}$ and $N_{ZH}\sim 10^{6}$ that indeed $\Delta^{NLO/LO}_{Z/ZH}$ can be $O(1)$. We may therefore expect competitive sensitivity at NLO on the $Z$ pole with respect to above-pole constraints at leading order for any Wilson coefficient that both have access to.
This argument also applies to new physics entering at leading order in both $Z$ pole and above-pole observables but with effects that grow with energy, as for the oblique $\hat{W}$ and $\hat{Y}$ parameters Maksymyk:1993zm ; Barbieri:2004qk . An energy enhancement at higher energies, for example in $WW$ measurements, can be compensated for by higher statistical accuracy on-pole if
@@
where we defined the improvement factor $\Delta^{LO/LO}$ to include energy-enhanced terms. For $WW$ energies $E_{WW}\sim 200$ GeV, $\epsilon_{Z}\ \sim 10^{-1}$, $\epsilon_{WW}\sim 1$, $N_{Z}\sim 10^{12}$ and $N_{WW}\sim 10^{8}$, we obtain $\Delta^{LO/LO}_{Z/WW}=O(1)$. We note that $N_{WW}\approx N_{\bar{f}f}\approx 10^{8}$, so the same argument applies for energy enhancement in the $e^{+}e^{-}\rightarrow\bar{f}f$ process. Thus, unlike at LEP (where $N_{Z}\sim 10^{7}$ and $N_{WW}\sim 10^{5}$), a Tera-Z factory is expected to be just as sensitive to the $\hat{W}$ and $\hat{Y}$ parameters on the $Z$ pole as in higher-energy $WW$ runs. Furthermore, it is important to combine all above-pole runs, as the number of fermion pairs produced falls off at roughly the same rate as the energy enhancement. This was recently shown explicitly by comparing $e^{+}e^{-}\rightarrow\bar{f}f$ above-pole processes with $Z$-pole data in Ref. Greljo:2024ytg .
In our analysis we will confirm this expectation, finding that many operators typically thought to be mainly constrained at leading order in the higher-energy $WW$, ZHZH, and $t\bar{t}$ runs can actually receive better or comparable constraints at NLO on the Z-pole. Moreover, including on-pole data eliminates directions of limited sensitivity in a global fit which can remove degeneracies and improve constraints when fitting to several correlated operators at a time, as is necessary in most realistic ultraviolet (UV) models.
@@
We indeed find this to be the case in several specific UV models that we study as examples. In the real scalar singlet and custodial quadruplet Durieux:2022hbu models, it is essential to take into account the correlated contributions to the Higgs self-coupling and Higgs self-energy when estimating the projected sensitivities in a global fit, even when there is a loop hierarchy between the two as for the weak quadruplet. Including these correlations weaken the bounds from above-pole measurements compared to considering only individual constraints; however, adding on-pole data then strengthens them further. For the real singlet scalar in particular, the on-pole information is crucial for potentially covering the entire region of parameter space in which the scalar can be a so-called “loryon” Banta:2021dek , thus closing the last remaining open window in being able to conclusively answer whether any other particles outside the SM get most of their mass from the Higgs Crawford:2024nun . We also consider the case of Weakly Interacting Massive Particles (WIMPs) and show that electroweak precision can significantly improve on the picture after HL-LHC for this elusive class of BSM extensions.
@@ 2 Setup and Observables
The measurements used in our analysis are classified as on-pole for $Z$- and $W$-pole observables, and as above-pole observables for everything else. For on-pole data, we use the following $Z$-pole observables,
@@
For these observables, we use the expressions calculated at NLO in the SMEFT with $G_{F}$, $m_{Z}$, $\alpha$} as input parameters in Ref. Bellafronte:2023amz . The relative uncertainties and projected error reductions for the $Z$- and $W$-pole observables are given in Table 1. The $W$ pole uncertainties are from Ref. Bernardi:2022hny . The $Z$ pole experimental uncertainties are taken from Ref. DeBlas:2019qco . We neglect theoretical uncertainties, assuming the “TH3” scenario of Ref. Blondel:2018mad . The sole exception to this is $\Gamma_{Z}$ Bernardi:2022hny ; Freitas:2019bre , which will likely remain theory limited with an error of $O(100)$ keV, which we use to determine the $\Gamma_{Z}$ improvement factor in Table 1.
@@
Table 1: Projected FCC-$e$ improvement for $Z$-pole observables and $W$ pole observables from Bernardi:2022hny ; DeBlas:2019qco . The $A_{\ell}^{**}$ are from lepton polarization and LR asymmetry measurements at SLC. For $A_{e}^{**}$, we use the projection in deBlas:2022ofj .
On the other hand, the above-pole processes considered here are
$O_{\text{above-pole}}=\left\{\sigma\left(e^{+}e^{-}\to W^{+}W^{-}\right),\,\sigma\left(e^{+}e^{-}\to ZH\right),\,\sigma\left(e^{+}e^{-}\to f\bar{f}\right)\,\right\}.$
For the $ZH$ production cross section, we use the results from Ref. Asteriadis:2024xts , which have been computed at NLO in the SMEFT. We use our own leading order (LO) calculation for the $WW$ process, given in Appendix A. For both calculations, the $G_{F}$, $m_{W}$, $m_{Z}$} input scheme is adopted. [^2] The choice of input parameter scheme modifies the predictions in two ways: it modifies the numerical value of the SM parameters used in the calculations and shifts the observable by a linear combination of $\mathcal{C}_{i}$. Using the results of Ref. Biekotter:2025nln , we have verified that the former is numerically very small. The latter effect is generally larger, but is confined to the coefficients $C_{HD},C_{HWB},[C_{Hl}^{(3)}]_{11},[C_{Hl}^{(3)}]_{22},[C_{ll}]_{1221}$ Biekotter:2023xle . These are constrained far more precisely at the Z-pole, so we do not expect the scheme difference has any impact on the sensitivity where these operators are concerned.
Regarding the $e^{+}e^{-}\to f\bar{f}$ process, we use the results of the recent dedicated flavor-tagging analysis for FCC-ee performed in Ref. Greljo:2024ytg . They use the cross-section ratio observables
@@
where the sum runs over all the light quarks. Our predictions are implemented following Appendix B of Ref. Greljo:2024ytg .
We first express all observables as functions of the Wilson coefficients at the scale of the observable, which we take to be $\mu=m_{Z}$ for the $Z/W$-pole and $\mu=\sqrt{s}=163,240,365$ GeV for observables above the pole. We then perform the RGE up to a high scale $\Lambda=1$ TeV to express all observables as a function of the high scale Wilson coefficients, which are then used to construct the likelihood. The fit is then performed via a simple $\chi^{2}$ procedure as in Ref. Allwicher:2023aql ; Allwicher:2023shc ; Stefanek:2024kds . The leading order $Z/W$-pole part of this likelihood has previously been applied and validated in Refs. Allwicher:2023aql ; Davighi:2023evx ; Allwicher:2023shc ; Stefanek:2024kds , and has also been cross-checked here with an independent implementation in fitmaker Ellis:2020unq . Finally, we note that the expected sensitivity for all observables has been appropriately rescaled to match the luminosity figures given in the FCC Feasibility Study Report Benedikt:2928193 .
Effective field theory framework
@@
where $Q_{i}$ are dimension-6 operators and $C_{i}$ are the dimensionless Wilson coefficients generated by heavy new physics at a scale $\Lambda$. We will take $\Lambda=1$ TeV unless otherwise stated.
In some cases, it will prove convenient to use the SILH basis Giudice:2007fh ; Contino:2013kra , such as to characterise pure gauge operators. To avoid conflict with Warsaw basis operators and Wilson coefficients, when working in the SILH basis we adopt the following notation
$\mathcal{L}_{\rm SMEFT}\supset\sum_{i}\frac{S_{i}}{\Lambda^{2}}\mathcal{O}_{i}^{\rm SILH}\,$, (4)
where the operators $\mathcal{O}_{i}^{\rm SILH}$ will always be explicitly defined in the text for convenience.
Finally, for more concise expressions we will sometimes employ dimensionful Wilson coefficients defined as $\mathcal{C}_{i}\equiv C_{i}/\Lambda^{2}$ and $\mathcal{S}_{i}\equiv S_{i}/\Lambda^{2}$.
3 Accuracy complements energy: an EFT analysis
The SMEFT language provides a systematic way of exploring modifications from heavy BSM physics that are correlated across different measurements at various energies (for some recent global fits, see e.g. Refs. Ellis:2020unq ; Bissmann:2020mfi ; Ethier:2021bye ; Grunwald:2023nli ; Garosi:2023yxg ; Allwicher:2023shc ; Bartocci:2023nvp ; Celada:2024mcf ; deBlas:2022ofj ; Bartocci:2024fmm ). Our approach is to focus on operators contributing to measurements above the $Z$ pole (denoted above-pole), some of which are energy enhanced, and that also enter the $Z$-pole at one higher loop order.
Figure 1: Projected sensitivity at FCC-ee of measurements on the $Z$ pole and above-pole at higher energy runs for Higgs, gauge, and four-fermion quark/lepton operators. The colour gradient denotes the bound on the scale $\Lambda$ in TeV for unit Wilson coefficient.
Fig. 1 summarises the complementarity between on-pole and above-pole constraints. The effects of heavy new physics are parametrised by the Wilson coefficients $\mathcal{C}_{i}\equiv C_{i}/\Lambda^{2}$ of dimension-6 operators $Q_{i}$ in the Warsaw basis of the SMEFT Grzadkowski:2010es , while the projected sensitivity at FCC-ee to the effective scale of new physics is represented by the colour-coded heatmap gradient. The operators shown are grouped by Higgs, gauge, and various four-fermion operators, and are typically thought to best be constrained at leading order above the pole at the $WW$, ZHZH, and $\bar{t}t$ higher-energy runs. Darker shading indicates a stronger constraint, and we see in many cases the on-pole constraint dominating or giving a similar order of magnitude sensitivity to the corresponding above-pole constraint. This is indeed the case for the $\hat{W}$ and $\hat{Y}$ oblique parameters that map on to the following operators in the SILH basis Giudice:2007fh ; Contino:2013kra
$\mathcal{L}_{\rm SMEFT}\supset-\frac{\mathcal{S}_{2W}}{2}(D^{\mu}W_{\mu\nu}^{I})(D_{\rho}W^{I\rho\nu})-\frac{\mathcal{S}_{2B}}{2}(\partial^{\mu}B_{\mu\nu})(\partial_{\rho}B^{\rho\nu})\,.$ (5)
For other operators the above-pole constraint can win due to energy enhancement effects, though the benefit of entering at leading order vanishes when considering four-fermion operators involving second and third generation fermions. Even when the sensitivity is worse, the complementarity of the on-pole constraints helps close flat directions in global fits and leads to better marginalised constraints or bounds on specific models.
We now illustrate in more detail how accuracy complements energy with Higgs sector modifications before considering pure gauge operators and then summarising the recent four-fermion results.
@@ 3.1 Higgs coupling modifications
Figure 4: New physics affecting the hVVhVV coupling in (a), NLO contribution of hVVhVV coupling modifications to $Z$-pole observables. Red squares indicate SMEFT contributions to the vertex.
A Higgs factory will be sensitive at leading order to new physics modifying the Higgs self-energy in the process $e^{+}e^{-}\to ZH$, as illustrated by the propagator correction in Fig. 2; we also see in Fig. 2 that it necessarily also enters at NLO in the oblique parameters on the $Z$ pole. Similarly, Fig. 3 shows that a vertex correction to the Higgs self-coupling will enter at NLO in $e^{+}e^{-}\to ZH$ and at NNLO on the $Z$ pole, while Fig. 4 shows a LO $hVV$ vertex correction modifying the $Z$ pole oblique parameter at NLO. As argued previously, we expect the suppression from an extra loop factor on the $Z$ pole to be compensated by increased statistics, such that the on-pole vs. above-pole improvement factor on constraints from all these Higgs coupling modifications could be $\Delta^{NLO/LO}_{Z/ZH}\sim\mathcal{O}(1)$ for both the ratios NLO/LO and NNLO/NLO.
@@
Figure 5: FCC-$e$ projected sensitivities at 68% CL to Higgs coupling modifications on and off the $Z/W$ pole. Left: SMEFT operators that modify the Higgs 2- and 3-point functions. The dashed region is obtained by linearizing Eq. 13. Right: SMEFT operators giving hVVhVV coupling modifications. The Wilson coefficients are renormalised at $\Lambda=1$ TeV.